### 06 Apr 2026 v1

@@ 2.2 Self-concordant convex-concave functions
    Further, if $\xi(w)\leq 0.1$, then $\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}<0.2$.
@@
The proofs of (i)(i)$(iv)$ follow directly from (Nemirovski, 1999, Proposition 2.3 and 5.1). The proof of $(v)$ is provided in Appendix A.2.
@@ 3 A minimax reformulation of the smoothing Newton method
In this section, we provide a minimax reformulation of the smoothing Newton method. We first review the classical SNM based on the smoothing CHKS function $\Phi$ for the SCP problem (1).
3.1 From smoothing CHKS function to the reduced SBAL function
Given any point $(x,s)\in\mathbb{E}\times\mathbb{E}$ and $\mu>0$, the classical SNM Engelke and Kanzow (2002); Kanzow and Nagel (2002); Liu et al. (2006) based on $\Phi$ for the SCP (1) (inexactly) solves the parameterized smooth equations
@@ 3.1 From smoothing CHKS function to the reduced SBAL function
Applying Newton’s method to the parameterized smooth equations (20) yields the linearized system:
@@
where $\mathcal{I}_{\mathbb{E}}:\mathbb{E}\to\mathbb{E}$ is an identity mapping. The classical SNM proceeds as follows. Starting from $(x^{(0)},s^{(0)},\lambda^{(0)})$, it computes the Newton direction given by (21), performs a line search along this direction at each iteration, and progressively updates the parameter $\mu$ toward zero.
To generalize the parameterized smooth equations, we derive an equivalent characterization of $\Phi$. Let $\phi(x)=-\ln\det(x)$ be the natural barrier of $\mathbb{K}$. Given a fixed $\rho\geq 1$, consider the following optimization problem:
$\min_{z\in\mathrm{int}\,(\mathbb{K})}\Big\{\mu\phi(z)+\langle s,z\rangle+\dfrac{\rho}{2}\|z-x\|^{2}\Big\}.$ (22)
The solution to (22) corresponds to the proximal mapping of the proper closed convex function $\mu\phi(\cdot)+\langle s,\cdot\rangle$ at $x$, which exists uniquely by (Beck, 2017, Theorem 6.3). We denote this solution by $z_{\rho}(x,s;\mu)$. Then
@@
and satisfies the optimality condition of (22):
@@
For brevity, we write $z_{\rho}$ instead of $z_{\rho}(x,s;\mu)$ whenever no confusion arises. The natural generalization of $\Phi$ follows as
$\Phi_{\rho}(x,s;\mu):=2(x-z_{\rho}(x,s;\mu)),\quad\forall x,s\in\mathbb{E},\ \mu>0,\ \text{and }\rho\geq 1.$ (24)
In particular, $\Phi(x,s;\mu)=\Phi_{1}(x,s;\mu)$. The parameterized smooth equations (20) generalize to
@@
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ D_{x}\Phi_{\rho}(x,s;\mu)&D_{s}\Phi_{\rho}(x,s;\mu)&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\mathcal{A}x-b\\ \mathcal{A}^{*}\lambda+s-c\\ \Phi_{\rho}(x,s;\mu)\end{pmatrix}$ (26)
The reformulation (25) extends the standard parameterized smooth equations system (20). Thus, we do not distinguish between the generalized SNM and the classical SNM, and simply refer to both as the SNM. In the following discussion, it suffices to focus on how to reformulate (25).
In line with the idea of building a connection between $\Phi_{\rho}$ and an optimization problem, we next relate (25) to a minimax problem. It is straightforward to verify that (25) is equivalent to
@@
$\mathcal{A}x=b,\;c-s-\rho(z_{\rho}-x)-\mathcal{A}^{*}\lambda=0,\;z_{\rho}-x=0,\;\mu\nabla\phi(z_{\rho})+s+\rho(z_{\rho}-x)=0.$ (28)
This system coincides exactly with the first-order optimality conditions of the following minimax problem, whose objective is the smoothing barrier augmented Lagrangian function $L_{\rho}$:
$\min\limits_{x,z}\max\limits_{s,\lambda}\ \left\{L_{\rho}(x,z,s,\lambda;\mu)\right\},$ (29)
where
$L_{\rho}(x,z,s,\lambda;\mu):=\langle{c},{x}\rangle+\mu\phi(z)-\langle\lambda,\mathcal{A}x-b\rangle+\langle s,z-x\rangle+\dfrac{\rho}{2}\left\|z-x\right\|^{2}.$ (30)
This reformulation reveals that the parameterized smooth equations (25) can be interpreted as the first-order optimality conditions of the minimax problem (29). Consequently, the corresponding theoretical properties of the SNM can be investigated via the SBAL function $L_{\rho}$. However, $L_{\rho}$ is degenerate in the $s$-variable, as $D_{ss}^{2}L_{\rho}(x,z,s,\lambda;\mu)=0$. To overcome this difficulty, we eliminate auxiliary variables and derive a reduced form of $L_{\rho}$.
Since $\mathcal{A}$ is surjective, the linear system $\mathcal{A}x=b$ admits a solution for any $b$. Fix an arbitrary feasible point $\bar{x}$ such that $\mathcal{A}\bar{x}=b$. Let $\hat{\mathbb{E}}$ be a finite-dimensional Euclidean space with ${\rm dim}\,\hat{\mathbb{E}}={\rm dim}\,{\rm ker}\,\mathcal{A}.$ Then, there exists an injective linear operator $\mathcal{B}:\hat{\mathbb{E}}\rightarrow\mathbb{E}$ such that
$\mathcal{A}\mathcal{B}=0,\ \quad\mathcal{B}^{*}\mathcal{B}=\mathcal{I}_{\hat{\mathbb{E}}}$,
@@
Restricting $L_{\rho}$ to the affine set $\{x\mid Ax=b\}$ eliminates the multiplier $\lambda$. We define $\eta_{\rho}(\cdot,\cdot;\mu):\hat{\mathbb{E}}\times\mathbb{E}\rightarrow\mathbb{R}$ by
$\eta_{\rho}(\hat{x},s;\mu)=\min\limits_{z}\left\{L_{\rho}(\bar{x}+\mathcal{B}\hat{x},z,s,\lambda;\mu)\right\},$ (31)
where the value is independent of $\lambda$ because $A(\bar{x}+\mathcal{B}\hat{x})=b$. Writing $x=\bar{x}+\mathcal{B}\hat{x}$, the unique minimizer of (31) is $z_{\rho}(x,s;\mu)$, and thus
@@
We refer to $\eta_{\rho}$ as the reduced SBAL function. The corresponding minimax problem is
$\min\limits_{\hat{x}\in\hat{\mathbb{E}}}\max\limits_{s\in\mathbb{E}}\left\{\eta_{\rho}(\hat{x},s;\mu)\right\}.$ (33)
Compared with $L_{\rho}$, the reduced SBAL function $\eta_{\rho}$ admits more favorable structural properties, as established in the following subsection. In the remainder of the paper, we focus exclusively on the minimax problem (33) and $\eta_{\rho}$.
3.2 Properties of the reduced SBAL function
In this subsection, we discuss the properties of the reduced SBAL function $\eta_{\rho}$. Recall that $z_{\rho}(x,s;\mu)$ satisfies the nonlinear equation:
@@ 3.2 Properties of the reduced SBAL function
$y_{\rho}(x,s;\mu)=s+\rho(z_{\rho}(x,s;\mu)-x).$
The variables $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$ satisfy the following properties.
@@
For any scalars $\mu>0$ and $\rho\geq 1$, the following statements are equivalent:
@@
Suppose that condition $(iii)$ holds. It follows from (34) that
$z_{\rho}(x,s;\mu)=\frac{\rho x-s+\left((s-\rho x)^{2}+4\rho\mu e\right)^{1/2}}{2\rho}.\\$ (36)
@@
$z_{\rho}(x,s;\mu)$ $=\frac{\rho x-s+\left(s^{2}-2\rho s\circ x+\rho^{2}x^{2}+4\rho\mu e\right)^{1/2}}{2\rho}=x,$ (37)
which establishes condition $(i)$.
Conversely, assume that condition $(i)$ holds. By (34) and the inclusions $z_{\rho},x\in\operatorname{int}\,(\mathbb{K})$,
@@
Combining (11) and (38) yields $s\in\operatorname{int}\,(\mathbb{K})$ and $x\circ s=\mu e$. This establishes condition $(iii)$ and completes the proof.
@@
For brevity, all derivatives with respect to $\mu$ are denoted by a prime. For example, $z_{\rho}^{\prime}(x,s;\mu)=D_{\mu}z_{\rho}(x,s;\mu)$. When no ambiguity arises, we also abbreviate $y_{\rho}(x,s;\mu)$ as $y_{\rho}$. The following theorem characterizes the derivatives of $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$.
@@
For any $\rho\geq 1$, the mappings $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$ are smooth with respect to $(x,s,\mu)$ on $\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++}$. Moreover, their partial derivatives with respect to $(x,s)$ are given by
@@
the Jacobian of (34) with respect to $z_{\rho}$ is nonsingular. Hence, the implicit function theorem implies $z\in C^{\infty}(\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++},\mathbb{E})$. By definition,
@@
By Theorem 3.1 and (26), the search direction generated by the SNM is equivalently written as the solution to the linear system
@@
The following proposition guarantees the uniqueness of the above search direction.
@@
For any point $(x,s,\lambda)$ and scalars $\mu>0$, $\rho\geq 1$, the Newton system (45) generated by the SNM admits a unique solution.
@@
$\Delta s=-\rho\mathcal{W}\Delta x-\rho\mathcal{H}(x-z_{\rho}).$ (46)
Substitute this expression into the remaining equations. It suffices to verify the uniqueness of the solution to
@@
The Schur complement of $\begin{pmatrix}-\rho\mathcal{W}&\mathcal{A}^{*}\\ \mathcal{A}&0\end{pmatrix}$ relative to $\rho\mathcal{W}$ is $\rho^{-1}\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}$ since $\rho\mathcal{W}$ is invertible. Therefore, the system (47) admits a unique solution. This completes the proof.
The next corollary provides explicit expressions for the first- and second-order derivatives of $\eta_{\rho}(\hat{x},s;\mu)$ with respect to $\hat{x}$ and $s$.
@@
For any $\rho\geq 1$, the reduced SBAL function $\eta_{\rho}(\hat{x},s;\mu)$ is smooth on $\hat{\mathbb{E}}\times\mathbb{E}\times\mathbb{R}_{++}$. Furthermore, for $x=\bar{x}+\mathcal{B}\hat{x}$,
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\mathcal{B}^{*}(c-y_{\rho}(x,s;\mu)),\,\nabla_{s}\eta_{\rho}(\hat{x},s;\mu)=z_{\rho}(x,s;\mu)-x,$ (48)
@@
$D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B},\quad D^{2}_{ss}\eta_{\rho}(\hat{x},s;\mu)=-\rho^{-1}\mathcal{H}^{-1}.$ (49)
@@
The smoothness of $\eta_{\rho}$ follows immediately from Theorem 3.1. By differentiating (31) with respect to $x$, we obtain from the equality $\hat{x}=\mathcal{B}^{*}(x-\bar{x})$ and the chain rule that
$\nabla_{\hat{x}}\eta_{\rho}(\hat{x},s;\mu)$ $=\mathcal{B}^{*}(\nabla_{x}\eta_{\rho}(\hat{x},s;\mu))$ (50)
@@
$\qquad+((D_{x}z_{\rho}(x,s;\mu)-\mathcal{I}_{\mathbb{E}})(s+\rho(z_{\rho}(x,s;\mu)-x))\big)$
$\overset{(i)}{=}\mathcal{B}^{*}\big(D_{x}z_{\rho}(x,s;\mu)(\mu\nabla\phi(z_{\rho}(x,s;\mu))+y_{\rho}(x,s;\mu))$
@@
$=\mathcal{B}^{*}(c-y_{\rho}(x,s;\mu)),$
where $(i)$ follows from the definition of $y_{\rho}(x,s;\mu)$. Further differentiating (50) with respect to $\hat{x}$ yields
$D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(\hat{x},s;\mu)=\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}.$ (51)
@@
For any $\mu>0$ and $\rho\geq 1$, the reduced SBAL function $\eta_{\rho}(\cdot,\cdot;\mu)$ is a nondegenerate $\mu$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. Furthermore, $\eta_{\rho}(\cdot,s;\mu)$ is nondegenerate $\mu$-self-concordant on $\hat{\mathbb{E}}$ for every $s\in\mathbb{E}$, and $\eta_{\rho}(\hat{x},\cdot;\mu)$ is nondegenerate $\mu$-self-concordant on $\mathbb{E}$ for every $\hat{x}\in\hat{\mathbb{E}}$.

### 21 May 2026 v2

@@ 2.2 $\alpha$-self-concordant convex-concave functions
    Furthermore, if $\xi(w)\leq 0.1$, then $\max\{\tilde{\delta}_{\hat{x}}(w),\tilde{\delta}_{s}(w)\}<0.2$.
@@
The proofs of (i)(i)$(iii)$ follow directly from (Nemirovski, 1999, Propositions 2.3 and 5.1). The proof of $(iv)$ is provided in Appendix A.2.
@@ 3 A minimax reformulation of the smoothing Newton method
In this section, we first reformulate the generalized parameterized smooth system as the first-order optimality conditions of a minimax problem. By eliminating the multiplier and the auxiliary variable, we obtain a reduced barrier augmented Lagrangian function $\eta_{\rho}$. We then show that $\eta_{\rho}$ is a nondegenerate $\mu$-self-concordant convex-concave function. Finally, we establish that the generalized parameterized smooth system is equivalent to the first-order optimality conditions of the minimax problem defined by $\eta_{\rho}$, and that the Newton system of SNM is equivalent to the corresponding Newton system for that minimax problem.
3.1 From the smoothing CHKS function to the reduced BAL function
We begin with the smoothing CHKS function
$\Phi(x,s;\mu)=x+s-\left((x-s)^{2}+4\mu e\right)^{1/2},$
and then introduce its generalized form $\Phi_{\rho}$.
Given any point $(x,s)\in\mathbb{E}\times\mathbb{E}$ and $\mu>0$, the classical SNM Engelke and Kanzow (2002); Kanzow and Nagel (2002); Liu et al. (2006) based on $\Phi$ for problem (1) (inexactly) solves the parameterized smooth system
@@ 3.1 From the smoothing CHKS function to the reduced BAL function
Applying Newton’s method to (18) yields the linearized system:
@@
where $\mathcal{I}_{\mathbb{E}}:\mathbb{E}\to\mathbb{E}$ denotes the identity operator on $\mathbb{E}$. The classical SNM proceeds as follows. Starting from $(x^{(0)},s^{(0)},\lambda^{(0)})$, it computes the Newton direction given by (19), performs a line search along this direction at each iteration, and progressively decreases $\mu$ toward zero.
To generalize the parameterized smooth system, we derive an equivalent characterization of $\Phi$. Let $\phi$ denote the natural barrier for $\mathbb{K}$. Given a fixed $\rho>0$, consider the following optimization problem:
$\min_{z\in\operatorname{int}(\mathbb{K})}\Big\{\mu\phi(z)+\langle s,z\rangle+\dfrac{\rho}{2}\|z-x\|^{2}\Big\}.$ (20)
Here $\rho$ serves as the proximal regularization parameter in the quadratic term $\frac{\rho}{2}\|z-x\|^{2}$. For every fixed $\rho>0$, the associated proximal problem (20) is well defined and admits a unique solution. The special choice $\rho=1$ recovers the classical smoothing CHKS function, whereas allowing arbitrary $\rho>0$ yields a family of smoothing CHKS-type functions.
The unique solution of (20) is the proximal point of the proper closed convex function $\mu\phi(\cdot)+\langle s,\cdot\rangle$ at $x$; see (Beck, 2017, Theorem 6.3). We denote this solution by $z_{\rho}(x,s;\mu)$. Then
@@
and satisfies the optimality condition
@@
Furthermore, $z_{\rho}(x,s;\mu)$ is a function of $(x,s;\mu)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++}$. For brevity, we write $z_{\rho}$ instead of $z_{\rho}(x,s;\mu)$ whenever no confusion arises. The natural generalization of $\Phi$ is then defined by
$\Phi_{\rho}(x,s;\mu):=2(x-z_{\rho}(x,s;\mu)),\quad\forall\,x,s\in\mathbb{E},\ \mu>0,\ \text{and }\rho>0.$ (22)
In particular, $\Phi(x,s;\mu)=\Phi_{1}(x,s;\mu)$. The parameterized smooth system (18) generalizes to
@@
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ D_{x}\Phi_{\rho}(x,s;\mu)&D_{s}\Phi_{\rho}(x,s;\mu)&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}\mathcal{A}x-b\\ \mathcal{A}^{*}\lambda+s-c\\ \Phi_{\rho}(x,s;\mu)\end{pmatrix}$. (24)
The reformulation (23) extends the parameterized smooth system (18), with the classical smoothing CHKS function recovered as the special case $\rho=1$. For simplicity, we continue to refer to the resulting method as an SNM.
We next show that (23) can be naturally related to a minimax problem, thereby further clarifying the optimization interpretation of $\Phi_{\rho}$. Indeed, (23) is equivalent to
@@
$\mathcal{A}x=b,\;c-s-\rho(z-x)-\mathcal{A}^{*}\lambda=0,\;z-x=0,\;\mu\nabla\phi(z)+s+\rho(z-x)=0.$
This system is precisely the system of first-order optimality conditions of the following minimax problem, whose objective is the barrier augmented Lagrangian (BAL) function $L_{\rho}$:
$\min\limits_{z\in\operatorname{int}(\mathbb{K}),x}\max\limits_{s,\lambda}\ \left\{L_{\rho}(x,z,s,\lambda;\mu)\right\},$ (26)
where $L_{\rho}(x,z,s,\lambda;\mu):=\langle{c},{x}\rangle+\mu\phi(z)-\langle\lambda,\mathcal{A}x-b\rangle+\langle s,z-x\rangle+\dfrac{\rho}{2}\left\|z-x\right\|^{2}.$ The idea underlying the BAL function can be traced back to Liu and Dai (2020), and was further developed in Liu et al. (2022, 2023); Zhang et al. (2023, 2024, 2026).
Reformulation (26) reveals that the parameterized smooth system (23) can be interpreted as the first-order optimality conditions of the minimax problem (26). This suggests studying the SNM through the BAL function $L_{\rho}$. However, $L_{\rho}$ is not suitable for direct self-concordant convex-concave analysis. Indeed, it is degenerate in both the multiplier $\lambda$ and the $s$-variable, since $D_{\lambda\lambda}^{2}L_{\rho}(x,z,s,\lambda;\mu)=0$ and $D_{ss}^{2}L_{\rho}(x,z,s,\lambda;\mu)=0.$ Therefore, we first restrict the formulation to the affine constraint $\mathcal{A}x=b$. Since $\mathcal{A}$ is surjective, the linear system $\mathcal{A}x=b$ admits a solution for the given $b$. Fix an arbitrary feasible point $\bar{x}$ such that $\mathcal{A}\bar{x}=b$. Let $\hat{\mathbb{E}}$ be a finite-dimensional Euclidean space such that ${\rm dim}\,\hat{\mathbb{E}}={\rm dim}\,{\rm ker}\,\mathcal{A}.$ Then there exists an injective linear operator $\mathcal{B}:\hat{\mathbb{E}}\rightarrow\mathbb{E}$ such that
$\mathcal{A}\mathcal{B}=0,\ \quad\mathcal{B}^{*}\mathcal{B}=\mathcal{I}_{\hat{\mathbb{E}}}$.
@@
Under this parametrization, the constraint term $\langle\lambda,\mathcal{A}x-b\rangle$ vanishes identically, and the multiplier $\lambda$ no longer appears. Thus, the degeneracy with respect to $\lambda$ is removed.
We then eliminate the auxiliary variable $z$. This reduction also removes the degeneracy in the $s$-variable and leads to a reduced formulation suitable for a self-concordant convex-concave analysis. Define $\eta_{\rho}(\cdot,\cdot;\mu):\hat{\mathbb{E}}\times\mathbb{E}\rightarrow\mathbb{R}$ by
$\eta_{\rho}(\hat{x},s;\mu)=\min\limits_{z\in\operatorname{int}(\mathbb{K})}\left\{L_{\rho}(\bar{x}+\mathcal{B}\hat{x},z,s,\lambda;\mu)\right\}.$ (27)
The value in $\eta_{\rho}(\hat{x},s;\mu)$ is independent of $\lambda$, because $\mathcal{A}(\bar{x}+\mathcal{B}\hat{x})=b$. Writing $x=\bar{x}+\mathcal{B}\hat{x}$, the unique minimizer in (27) is $z_{\rho}(x,s;\mu)$. Hence
@@
We call $\eta_{\rho}$ the reduced BAL function. In the sequel, we write
$w:=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$
for the reduced minimax variable whenever no confusion can arise. The resulting reduced minimax problem is
$\min_{\hat{x}\in\hat{\mathbb{E}}}\max_{s\in\mathbb{E}}\left\{\eta_{\rho}(w;\mu)\right\}.$ (29)
Compared with the original BAL function $L_{\rho}$, the reduced function $\eta_{\rho}$ removes the affine multiplier degeneracy and, after eliminating $z$, has a nondegenerate curvature structure in both the minimization and maximization variables. These structural properties are established in the next subsection. In the remainder of the paper, we work with the reduced minimax problem (29) and the function $\eta_{\rho}$.
3.2 Properties of the reduced BAL function
In this subsection, we discuss the properties of the reduced BAL function $\eta_{\rho}$. Recall that $z_{\rho}(x,s;\mu)$ satisfies
@@ 3.2 Properties of the reduced BAL function
$y_{\rho}(x,s;\mu):=s+\rho(z_{\rho}(x,s;\mu)-x).$ (31)
The variable $y_{\rho}(x,s;\mu)$ is also a function of $(x,s;\mu)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++}$. Furthermore, both $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$ satisfy the following properties.
@@
For any scalars $\mu>0$ and $\rho>0$, the following statements are equivalent:
@@
Suppose that $(iii)$ holds. By (30),
$z_{\rho}(x,s;\mu)=\frac{\rho x-s+\left((s-\rho x)^{2}+4\rho\mu e\right)^{1/2}}{2\rho}.\\$
@@
$z_{\rho}(x,s;\mu)$ $=\frac{\rho x-s+\left(s^{2}-2\rho s\circ x+\rho^{2}x^{2}+4\rho\mu e\right)^{1/2}}{2\rho}=x,$
which establishes $(i)$.
Conversely, assume that $(i)$ holds. By (30) and the inclusions $z_{\rho},x\in\operatorname{int}(\mathbb{K})$,
@@
Combining (10) and (33) yields $s\in\operatorname{int}(\mathbb{K})$ and $x\circ s=\mu e$. This establishes $(iii)$ and completes the proof.
@@
For brevity, all derivatives with respect to $\mu$ are denoted by a prime. For example, $z_{\rho}^{\prime}(x,s;\mu):=D_{\mu}z_{\rho}(x,s;\mu)$. When no ambiguity arises, we also abbreviate $y_{\rho}(x,s;\mu)$ as $y_{\rho}$. The following theorem characterizes the derivatives of $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$.
@@
For any $\rho>0$, the functions $z_{\rho}(x,s;\mu)$ and $y_{\rho}(x,s;\mu)$ are smooth with respect to $(x,s,\mu)$ on $\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++}$. Furthermore, their partial derivatives with respect to $(x,s)$ are given by
@@
the derivative of the mapping
$F_{\rho}(x,z,s;\mu)=\mu\nabla\phi(z)+s+\rho(z-x)\in C^{\infty}(\mathbb{E}\times\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++},\mathbb{E})$
with respect to $z$ at $z=z_{\rho}(x,s;\mu)$ is nonsingular. Hence, the implicit function theorem implies $z_{\rho}\in C^{\infty}(\mathbb{E}\times\mathbb{E}\times\mathbb{R}_{++},\mathbb{E})$. By definition,
@@
By Theorem 3.1 and (24), the search direction generated by the SNM is equivalently written as the solution of the linear system
@@
The following proposition guarantees the uniqueness of this search direction.
@@
For any point $(x,s,\lambda)\in\mathbb{E}\times\mathbb{E}\times\mathbb{R}^{m}$ and scalars $\mu>0$, $\rho>0$, the Newton system (40) generated by the SNM admits a unique solution.
@@
$\Delta s=-\rho\mathcal{W}\Delta x-\rho\mathcal{H}(x-z_{\rho}).$
Substituting this expression into the remaining equations, it suffices to verify the uniqueness of the solution to
@@
Since $\rho\mathcal{W}$ is invertible, the Schur complement of $\begin{pmatrix}-\rho\mathcal{W}&\mathcal{A}^{*}\\ \mathcal{A}&0\end{pmatrix}$ relative to $\rho\mathcal{W}$ is $\rho^{-1}\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}$. Moreover, $\mathcal{W}^{-1}\succ 0$ and $\mathcal{A}$ is surjective. Therefore, $\rho^{-1}\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}$ is invertible and the system (41) admits a unique solution.
The next corollary provides explicit formulas for the first- and second-order derivatives of $\eta_{\rho}(w;\mu)$ with respect to $\hat{x}$ and $s$.
@@
For any $\rho>0$, the reduced BAL function $\eta_{\rho}(w;\mu)$ is smooth on $\hat{\mathbb{E}}\times\mathbb{E}\times\mathbb{R}_{++}$. Furthermore, for $x=\bar{x}+\mathcal{B}\hat{x}$,
$\nabla_{\hat{x}}\eta_{\rho}(w;\mu)=\mathcal{B}^{*}(c-y_{\rho}(x,s;\mu)),\,\nabla_{s}\eta_{\rho}(w;\mu)=z_{\rho}(x,s;\mu)-x,$ (42)
@@
$D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(w;\mu)$ $=\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B},\quad D^{2}_{ss}\eta_{\rho}(w;\mu)=-\rho^{-1}\mathcal{H}^{-1},$ (43)
$D^{2}_{\hat{x}s}\eta_{\rho}(w;\mu)$ $=-\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W},\quad D^{2}_{s\hat{x}}\eta_{\rho}(w;\mu)=-\mathcal{H}^{-1}\mathcal{W}\mathcal{B}.$
@@
The smoothness of $\eta_{\rho}$ follows immediately from Theorem 3.1. By differentiating (27) with respect to $\hat{x}$ and using the chain rule, we obtain
$\nabla_{\hat{x}}\eta_{\rho}(w;\mu)$ $=\mathcal{B}^{*}(\nabla_{x}\eta_{\rho}(w;\mu))$
@@
$\qquad+(D_{x}z_{\rho}(x,s;\mu)-\mathcal{I}_{\mathbb{E}})\left(s+\rho(z_{\rho}(x,s;\mu)-x)\right)\big).$
By (30) and (31), we have
$\nabla_{\hat{x}}\eta_{\rho}(w;\mu)$ $=\mathcal{B}^{*}\big(D_{x}z_{\rho}(x,s;\mu)(\mu\nabla\phi(z_{\rho}(x,s;\mu))+y_{\rho}(x,s;\mu))$ (44)
@@
$=\mathcal{B}^{*}(c-y_{\rho}(x,s;\mu)).$
Further differentiating (44) with respect to $\hat{x}$ yields
$D^{2}_{\hat{x}\hat{x}}\eta_{\rho}(w;\mu)=\rho\mathcal{B}^{*}\mathcal{H}^{-1}\mathcal{W}\mathcal{B}.$
@@
Corollary 1 implies that $\eta_{\rho}(w;\mu)$ is convex in $\hat{x}$ and concave in $s$. The following theorem further shows that $\eta_{\rho}(w;\mu)$ is a nondegenerate $\mu$-self-concordant convex–concave function.
@@
For any $\mu>0$ and $\rho>0$, the reduced BAL function $\eta_{\rho}(\cdot,\cdot;\mu)$ is a nondegenerate $\mu$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. Furthermore, $\eta_{\rho}(\cdot,s;\mu)$ is nondegenerate $\mu$-self-concordant on $\hat{\mathbb{E}}$ for every $s\in\mathbb{E}$, and $\eta_{\rho}(\hat{x},\cdot;\mu)$ is nondegenerate $\mu$-self-concordant on $\mathbb{E}$ for every $\hat{x}\in\hat{\mathbb{E}}$.